OpenStudy (honeymonster):

please help me solve this linear equation..find the value of x..i'll post the equation in replies..i

OpenStudy (honeymonster):

\[\frac{ x }{ 2 }+\frac{ x }{ 3 }=1\]

hartnn (hartnn):

first multiply 6 on both sides and tell me what u get?

OpenStudy (anonymous):

reasoning for multiplying by 6 first determine the GCD-greatest common denominator of the 2 fractions the GCD turns out to be 6 convert the fractions so they both have a denominator of 6 then combine the 2 fractions because they both have denominator of 6 now because you want to get rid of the denominator, you multiply by 6 to both sides of the equation

OpenStudy (honeymonster):

i dont understand :/

OpenStudy (anonymous):

do you understand what a GCD is ?

OpenStudy (honeymonster):

not really :/ sorry :/

OpenStudy (anonymous):

you do know what a denominator is right?

OpenStudy (honeymonster):

yes..

OpenStudy (anonymous):

hmm i might've meant LCD not GCD w/e anyways say we have 1/2 and 1/3 we want to find the lowest denominator that is a multiply of both 2 and 3 12 is a multiply of 2 and 3, but its not the lowest 6 is also a multiply of 2 and 3 and it turns out to be the lowest denominator

OpenStudy (anonymous):

now we want to convert 1/2 and 1/3 into fractions with denominators of 6 \[\frac{1}{2}*\frac{3}{3}=\frac{3}{6}\] \[\frac{1}{3}*\frac{2}{2}=\frac{2}{6}\] by doing so, we can now add the 2 fractions together \[\frac{3}{6}+\frac{2}{6}=\frac{3+2}{6}=\frac{5}{6}\]

OpenStudy (anonymous):

does this make sense?

OpenStudy (honeymonster):

yes that does make sense..

OpenStudy (anonymous):

we can apply this to your question \[\frac{x}{2}=\frac{1}{2}x=\frac{3}{3}*\frac{1}{2}x=\frac{3}{6}x\]

OpenStudy (anonymous):

add up the 2 fractions, then solve for x

OpenStudy (honeymonster):

ok thank you for your help :)

OpenStudy (anonymous):

no prob